# Baryon Pasting the Uchuu Lightcone Simulation

### v1

@@ 1 Introduction
Accurate estimation of the mass-observable scaling relations is fundamental to derive reliable constraints on key cosmological parameters in cluster abundance cosmology (Pratt et al., 2019). These scaling relations connect the halo mass to observable properties, such as X-ray luminosity or the thermal Sunyaev–Zeldovich (tSZ) signal at microwave wavelengths, enabling indirect mass estimation in cosmological surveys (e.g., Vikhlinin et al., 2009; Mantz et al., 2010; Benson et al., 2013; Bocquet et al., 2019). However, these relations often fail to capture the full complexity of physical systems because they do not account for the scatter around the mean relations (e.g., Mantz et al., 2016; Sereno et al., 2020). While scatter provides direct information on the structure and evolution of these systems (Farahi et al., 2019), failing to account for scatter biases in the interpretation of observational data (Farahi et al., 2019; Costanzi et al., 2019; Zhang et al., 2024), thereby affecting the precision estimates of crucial cosmological parameters such as the matter density ($\Omega_{m}$) and the amplitude of matter fluctuations ($\sigma_{8}$).
Scatter in scaling relations can be categorized into two primary sources: intrinsic and extrinsic. Intrinsic scatter arises from internal physical processes within the halo, such as baryonic physics such as feedback from supernovae (SNe) and active galactic nuclei (AGN), radiative gas cooling, and star formation, all of which influence the thermodynamics of the halo gas (e.g., Farahi et al., 2018; Truong et al., 2018; Pop et al., 2022; Yang et al., 2022). Extrinsic scatter, on the other hand, originates from external factors such as the non-spherical (triaxial) nature of halos and projection effects. Similarly to dark matter (DM) distributions, the gas distributions in clusters and groups are typically triaxial (Lau et al., 2011; Mulroy et al., 2019; Kim et al., 2024), and the observed signals can be affected by the orientation of the halo’s major axis relative to the line of sight (e.g., Limousin et al., 2013). Understanding and quantifying both intrinsic and extrinsic scatter is essential to improve the accuracy of mass estimates and minimizing biases in cosmological inferences.
@@
To address the challenges of modeling halo gas for large-scale multiwavelength galaxy surveys, we have developed Baryon Pasting (BP), a code specifically designed to paint gas onto DM halos efficiently and fast, with physically motivated models. BP provides physically interpretable modeling of gas in DM halos, which is essential for astrophysical and cosmological inferences from observations. It enables rapid exploration of how feedback physics and non-thermal pressure support affect the tSZ power spectrum (Shaw et al., 2010), how cool-dense cores affect optical depth measurements of clusters and groups in kinetic SZ observations (Flender et al., 2017), and how to break the degeneracies between astrophysical and cosmological parameters in cross-correlations of X-ray, tSZ, and lensing measurements of galaxy clusters and groups (Shirasaki et al., 2020). Painting gas on DM particles in the simulation allows us to model gas observables in cosmic web filaments (Osato & Nagai, 2023). Unlike purely empirical or machine learning-based methods, BP maintains physical interpretability while remaining computationally efficient, making it well-suited for the extensive parameter space exploration required by upcoming surveys.
@@ 2.1 Halo Gas Profiles
Here we briefly review the salient features of the core BP gas model, which are described in detail in Shaw et al. (2010); Flender et al. (2017); Osato & Nagai (2023). The basic assumption of the BP model is that the DM halo follows the profile Navarro–Frenk–White (NFW) profile (Navarro et al., 1996),
@@
and ${\Gamma=1+1/n}=1.2$ and $n=5$ are the polytropic exponent and the polytropic index respectively, whose values are set to match those in cosmological hydrodynamical simulations (Komatsu & Seljak, 2001; Ostriker et al., 2005; Shaw et al., 2010; Battaglia et al., 2012). The values are also consistent with recent observations of the polytropic index (Ghirardini et al., 2019).
@@
$\Gamma=\begin{cases}1.2&r_{\rm break}/R_{500c}\geq x_{\rm break},\\ \Gamma_{0}(1+z)^{\beta}&r_{\rm break}/R_{500c}<x_{\rm break},\end{cases}$ (6)
@@
where $R_{\rm 200m}$ is the spherical over-density radius with respect to 200 times the mean matter density of the universe. The parameters are calibrated to be $A_{\rm nt}=0.452$, $B_{\rm nt}=0.841$, $C_{\rm nt}=1.628$. The $R_{\rm 200m}$ scaling ensures halo redshift and mass independence at the cluster scales $M_{500c}\geq 3\times 10^{14}h^{-1}M_{\odot}$ over which this relation is calibrated with the Omega500 cosmological simulation (Nelson et al., 2014a). The thermal pressure $P_{\rm th}$ is obtained by multiplying $P_{\rm tot}$ with $(1-f_{\rm nt})$.
Table 1 summarizes the parameters of the BP gas model, their physical meanings, and default values.
@@ 2.2 Calculating X-ray and thermal SZ Observables
The Compton-$y$ profile of a given halo is directly derived from the thermal pressure profile:
@@ 2.3 Triaxial Gas Halo and its 2D Projection
where $\mu_{14}\equiv\log_{10}(M_{\rm vir}/10^{14}M_{\odot})$ and $\zeta\equiv\log_{10}(1+z)$. Note that these relations are calibrated on the TNG300 simulations. In principle, they depend on the underlying baryon physics (Kazantzidis et al., 2004), which varies between different simulations.
@@
where $z_{\star}=\sqrt{f}\left(z^{\prime}+g/(2f)\right)$, $\zeta=h-{g^{2}}/{(4f)}$, and we set $r_{\rm max}=\sqrt{f}\left(5R_{500c}+g/(2f)\right)$.
@@ 3.1 Uchuu Lightcone
The Uchuu Lightcone contains halos with a mass range of $M_{500c}\geq 10^{13}M_{\odot}$ and a redshift range of $0\leq z\leq 2$, with a total of $75,159,192$ halos. The lightcone catalog also contains information on the ratios of the halo axis $c/a$ and $b/a$, and the orientation of the major axis, derived from the ROCKSTAR halo catalog.
@@ 3.2 Generation of the Baryon Pasted Maps
with a total of 8 maps (4 X-ray Surface Brightness, 4 Compton-$y$). Details of the triaxial halo projection and intrinsic scatter modeling can be found in Sections 2.3 and 2.4, respectively. Note that no foreground or background noise is applied to these maps. Figure 4 shows the X-ray and Compton-$y$ maps with triaxiality but no intrinsic scatter (the ‘tri’ maps) in the gas profiles. Note that maps with triaxial halos with intrinsic scatter (‘tri+var’) overestimate the level of total scatter, since intrinsic scatter that we derived empirically from cosmological simulations does include the contribution from triaxiality. The scatter derived from this map thus provides an upper limit to the total amount of scatter expected from both intrinsic scatter and triaxiality.
Figure 5: Left panel: X-ray luminosity-Mass scaling relation with the X-ray luminosity $L_{X}$ extracted from circular apertures with radii $R_{500c}$ in cluster-size halos $M_{500c}\geq 5\times 10^{13}M_{\odot}$ at $z<0.5$ from the X-ray Surface Brightness Uchuu BP Map, generated with halo triaxiality (the ‘tri’ map). The $E(z)^{-2}=H(z=0)/H(z)$ factor accounts for the self-similar redshift evolution of $L_{X}$. The color indicates the magnitude of $\cos(\theta)$, where $\theta$ is the angle between the major axis of the halo and the line-of-sight. for the halos in the baryon pasted maps. Right panel: the ratio between the $L_{X}$ in the ‘tri’ map, to that extracted from the spherical ‘sph’ map.
Figure 6: This plot shows the ratio between the $L_{X}$ extracted from the halos in the triaxial BP map (‘tri’), to that extracted from the spherical BP map (‘sph’), as a function of the orientation magnitude $\cos(\theta)$. The color indicates the minor-to-major axis ratio $c/a$.
Figure 7: This plot shows the distribution of the $L_{X}$ extracted from the halos in the triaxial XSB map (‘tri’, solid lines), in the XSB map with spherical halos plus intrinsic scatter (‘sph+var’, dashed lines), and XSB map with triaxial halos with intrinsic scatter (‘tri+var’, dot-dasahed lines), to that extracted from the spherical BP XSB map (‘sph’), in different halo mass bins.
@@ 4.1 Extrinsic Scatter in X-ray Luminosity–Halo Mass Scaling Relation
Excluding halo core regions can significantly reduce intrinsic scatter in X-ray luminosity. By omitting pixels within a circular aperture of $0.2R_{500c}$ around each halo’s center, the intrinsic scatter decreases from $29\%$ to $22\%$ for groups and from $20\%$ to $11\%$ for massive clusters. However, the scatter due to triaxiality remains unchanged after core excision, maintaining at approximately $12\%$ at the group scale and $7\%$ at the cluster scale, constituting nearly half of the total scatter. These results indicate that the halo outskirts are the primary contributors to triaxial scatter.
@@ 4.2 Bias in X-ray and Thermal SZ Power Spectra due to Intrinsic Scatter in ICM Profiles
Figure 8: Plots showing the effect of intrinsic scatter of pressure and density profiles on the X-ray and tSZ auto- and cross-angular power spectra. The left, middle, and right panels show the X-ray auto-power spectra, X-ray/tSZ cross-power spectra, and tSZ auto-power spectra, measured from the Uchuu BP maps respectively. In each panel, the blue and orange lines represent the power spectra with and without intrinsic scatter in the halo profiles. The bottom sub-panels show the ratios of the power spectra with intrinsic scatter in the halo profiles, to those without. Note that the errorbars in each of the power spectrum are much smaller than the difference between the spectra. Also note that the tSZ auto-power spectrum is unit-less.
@@
The angular power spectrum of clusters is dominated by the one-halo term $C^{\rm 1h}$ at most scales of interest ($\ell>10$), so we will ignore the two-halo term. The observables $\cal A$ and $\cal B$ for a halo in a given mass and redshift bin are expected to deviate from their expected value in that bin
${\cal A}$ $=$ $\overline{\cal A}+\delta{\cal A},$ (33)
${\cal B}$ $=$ ¯$\overline{\cal B}+\delta{\cal B}$
where the overline indicates the expectation value and the $\delta$ terms indicate the deviation value which are treated as random variables. The $C^{\rm 1h}$ term then becomes
@@
×$\times\left(\langle\overline{{\cal A}}^{*}_{\ell}(M,z)\overline{{\cal B}}_{\ell}(M,z)\rangle+\langle\delta{\cal A}^{*}_{\ell}(M,z)\delta{\cal B}_{\ell}(M,z)\rangle\right),$
where * superscript indicates the complex conjugate. Note that the terms ${\cal A}^{*}\delta{\cal B}$ and ${\cal B}\delta{\cal A}^{*}$ become zero when they are averaged. If there is a correlation between the variations of the two observables, expressed as $\langle\delta{\cal A}^{*}\delta{\cal B}\rangle\neq 0$, the implications are as follows: if $\langle\delta{\cal A}^{*}\delta{\cal B}\rangle>0$, this leads to an enhanced power spectrum, while $\langle\delta{\cal A}^{*}\delta{\cal B}\rangle<0$ results in a reduction of power. For the auto-power spectrum, $\langle\delta{\cal A}^{*}\delta{\cal A}\rangle=\langle|\delta{\cal A}|^{2}\rangle\geq 0$, indicating that any noticeable fluctuations in the observable will invariably increase the clustering power. In the X-ray/tSZ cross-power spectrum, we expect an increase in power, since X-ray surface brightness, which scales with the square of the gas density, correlates positively with thermal pressure contributing to the tSZ signal.
Interpretation of the power spectrum that uses the halo model usually overlooks the intrinsic scatter in halo gas profiles. As a result, the halo model approach typically underestimates the actual amplitude of the power spectrum. Thus, cosmological parameters, such as $S_{8}$, inferred through statistical inference based on the halo model, are often overestimated to account for the additional power from intrinsic scatter.

### v2

@@ 1 Introduction
Accurate estimation of the mass-observable scaling relations is fundamental to derive reliable constraints on key cosmological parameters in cluster abundance cosmology (Pratt et al., 2019). These scaling relations connect the halo mass to observable properties, such as X-ray luminosity or the thermal Sunyaev–Zeldovich (tSZ) signal at microwave wavelengths, enabling indirect mass estimation in cosmological surveys (e.g., Vikhlinin et al., 2009; Mantz et al., 2010; Benson et al., 2013; Bocquet et al., 2019). However, these relations often fail to capture the full complexity of physical systems because they do not account for the scatter around the mean relations (e.g., Mantz et al., 2016; Sereno et al., 2020). While scatter provides direct information on the structure and evolution of these systems (Farahi et al., 2019a), failing to account for scatter biases in the interpretation of observational data (Farahi et al., 2019b; Costanzi et al., 2019; Zhang et al., 2024), thereby affecting the precision estimates of crucial cosmological parameters such as the matter density ($\Omega_{m}$) and the amplitude of matter fluctuations ($\sigma_{8}$).
Scatter in scaling relations can be categorized into two primary sources: intrinsic and extrinsic. Intrinsic scatter arises from internal physical processes within the halo, such as baryonic physics such as feedback from supernovae (SNe) and active galactic nuclei (AGN), radiative gas cooling, and star formation, all of which influence the thermodynamics of the halo gas (e.g., Farahi et al., 2018; Truong et al., 2018; Pop et al., 2022; Yang et al., 2022). Extrinsic scatter, on the other hand, originates from external factors such as the non-spherical (triaxial) nature of halos and projection effects. Similarly to dark matter (DM) distributions, the gas distributions in clusters and groups are typically triaxial (Lau et al., 2011; Mulroy et al., 2019; Kim et al., 2024), and the observed signals can be affected by the orientation of the halo’s major axis relative to the line of sight (e.g., Battaglia et al., 2012a; Limousin et al., 2013). Understanding and quantifying both intrinsic and extrinsic scatter is essential to improve the accuracy of mass estimates and minimizing biases in cosmological inferences.
@@
To address the challenges of modeling halo gas for large-scale multiwavelength galaxy surveys, we have developed Baryon Pasting (BP), a code specifically designed to paint gas onto DM halos efficiently and fast, with physically motivated models. BP provides physically interpretable modeling of gas in DM halos, which is essential for astrophysical and cosmological inferences from observations. It enables rapid exploration of how feedback physics (Ostriker et al., 2005; Bode et al., 2009; Trac et al., 2011) and non-thermal pressure support affect the tSZ power spectrum (Shaw et al., 2010), how cool-dense cores affect optical depth measurements of clusters and groups in kinetic SZ observations (Flender et al., 2017), and how to break the degeneracies between astrophysical and cosmological parameters in cross-correlations of X-ray, tSZ, and lensing measurements of galaxy clusters and groups (Shirasaki et al., 2020). Painting gas on DM particles in the simulation allows us to model gas observables in cosmic web filaments (Osato & Nagai, 2023). Unlike purely empirical or machine learning-based methods, BP maintains physical interpretability while remaining computationally efficient, making it well-suited for the extensive parameter space exploration required by upcoming surveys.
@@ 2.1 Halo Gas Profiles
Here we briefly review the salient features of the core BP gas model. The BP gas model builds on a simple feedback model by Ostriker et al. (2005), which was then extended to include additional physical effects, such as non-thermal pressure (Shaw et al., 2010) and cool-cores (Flender et al., 2017). We refer the reader to these papers for details. The basic assumption of the BP model is that the DM halo follows the profile Navarro–Frenk–White (NFW) profile (Navarro et al., 1996),
@@
and ${\Gamma=1+1/n}=1.2$ and $n=5$ are the polytropic exponent and the polytropic index respectively, whose values are set to match those in cosmological hydrodynamical simulations (Komatsu & Seljak, 2001; Ostriker et al., 2005; Shaw et al., 2010; Battaglia et al., 2012a). The values are also consistent with recent observations of the polytropic index (Ghirardini et al., 2019).
@@
$\Gamma=\begin{cases}1.2&r_{\rm break}/R_{500c}\geq x_{\rm break},\\ \Gamma_{0}(1+z)^{\beta_{g}}&r_{\rm break}/R_{500c}<x_{\rm break},\end{cases}$ (6)
@@
where $R_{\rm 200m}$ is the spherical over-density radius with respect to 200 times the mean matter density of the universe. The parameters are calibrated to be $A_{\rm nt}=0.452$, $B_{\rm nt}=0.841$, $C_{\rm nt}=1.628$. The $R_{\rm 200m}$ scaling ensures halo redshift and mass independence at the cluster scales $M_{500c}\geq 3\times 10^{14}h^{-1}M_{\odot}$ over which this relation is calibrated with the Omega500 cosmological simulation (Nelson et al., 2014a). The thermal pressure $P_{\rm th}$ is obtained by multiplying $P_{\rm tot}$ with $(1-f_{\rm nt})$. Note that in our model, instead of modeling the thermal pressure $P_{\rm th}$ directly, $P_{\rm th}$ is derived from the total pressure $P_{\rm tot}$ and the non-thermal pressure fraction $f_{\rm nt}$. Physically, this is because $P_{\rm tot}$ can be better described by a polytropic equation of state than that of $P_{\rm th}$. Specifically, $P_{\rm tot}$ results in a constant polytropic index across a wide range of gas density, while $P_{\rm th}$ requires a density-dependent $\Gamma$ (see Figure 1 in Shaw et al. 2010).
Table 1 summarizes the parameters of the BP gas model, their physical meanings, and default values. The BP model is flexible for describing gas profiles in cosmological simulations and observations. In Appendix A, we show the comparison of the density and pressure profiles between the BP model and those of the IllustrisTNG300 simulations and observations.
@@ 2.2 Calculating X-ray and thermal SZ Observables
The Compton-$y$ profile of a given halo is directly derived from the pressure profile of thermal electrons:
@@ 2.3 Triaxial Gas Halo and its 2D Projection
where $\mu_{14}\equiv\log_{10}(M_{\rm vir}/10^{14}M_{\odot})$ and $\zeta\equiv\log_{10}(1+z)$. Figure 2 shows the relations between DM and gas triaxial ratios and their best-fits. Note that these relations are calibrated on the TNG300 simulations. In principle, they depend on the underlying baryon physics (Kazantzidis et al., 2004), which varies between different simulations.
@@
where $z_{\star}=\sqrt{f}\left(z^{\prime}+g/(2f)\right)$, $\zeta=h-{g^{2}}/{(4f)}$, and we set $r_{\rm max}=\sqrt{f}\left(5R_{500c}+g/(2f)\right)$, corresponding to the outer accretion shock radius of the gaseous halo at $\sim 5R_{500c}$ (Aung et al., 2021).
@@ 3.1 Uchuu Lightcone
The Uchuu Lightcone contains halos with a mass range of $M_{500c}\geq 10^{13}M_{\odot}$ and a redshift range of $0\leq z\leq 2$, with a total of $75,159,192$ halos. The lightcone catalog also contains information on halo concentrations, ratios of the halo axis $c/a$ and $b/a$, and the orientation of the major axis, derived from the ROCKSTAR halo catalog.
@@ 3.2 Generation of the Baryon Pasted Maps
with a total of 8 maps (4 X-ray Surface Brightness, 4 Compton-$y$). Details of the triaxial halo projection and intrinsic scatter modeling can be found in Sections 2.3 and 2.4, respectively. Note that no foreground or background noise is applied to these maps. Figure 4 shows the X-ray and Compton-$y$ maps with triaxiality but no intrinsic scatter (the ‘tri’ maps) in the gas profiles. Note that maps with triaxial halos with intrinsic scatter (‘tri+var’) overestimate the level of total scatter, since intrinsic scatter that we derived empirically from cosmological simulations does include the contribution from triaxiality. The scatter derived from this map thus provides an upper limit to the total amount of scatter expected from both intrinsic scatter and triaxiality. Note that for these sets of maps, we do not use the halo concentration provided by ROCKSTAR to model the gas profiles. This is because halo concentration can contribute to extra scatter in the profile, and is also correlated with halo triaxiality (Lau et al., 2021). To avoid double-counting the scatter, we fix the halo concentration to be the same for halos with the same mass and redshift using the fitting function from Diemer & Kravtsov (2015).
Figure 5: The Left panel shows X-ray luminosity-Mass scaling relation with the X-ray luminosity $L_{X}$ extracted from circular apertures with radii $R_{500c}$ in cluster-size halos $M_{500c}\geq 5\times 10^{13}M_{\odot}$ at $z<0.5$ from the X-ray Surface Brightness Uchuu BP Map, generated with halo triaxiality (the ‘tri’ map). The $E(z)^{-2}=H(z=0)/H(z)$ factor accounts for the self-similar redshift evolution of $L_{X}$. The color indicates the magnitude of $\cos(\theta)$, where $\theta$ is the angle between the major axis of the halo and the line-of-sight. Halos with high values of $\cos(\theta)$, i.e. halos with major axes more aligned with the line-of-sight, drives up the scatter in $L_{X}$. This is more evident in the right pane where we show the ratio between the $L_{X}$ in the ‘tri’ map, to that extracted from the spherical ‘sph’ map. Triaxial halos with higher (lower) $\cos(\theta)$ shows larger (smaller) $L_{X}$ values compared to their spherical counterparts.
Figure 6: This plot shows the ratio between the $L_{X}$ extracted from the halos in the triaxial BP map (‘tri’), to that extracted from the spherical BP map (‘sph’), as a function of the orientation magnitude $\cos(\theta)$. The color indicates the minor-to-major axis ratio $c/a$. It shows that the scatter in $L_{X}$ due to triaxiality is driven mostly by elongated halos with $c/a\lesssim 0.5$. They have lower projected $L_{X}$ when their major axis is lying closer to the plane of sky (e.g., $|\cos\theta|<0.8$ ) and drive the projected $L_{X}$ values high when their major axes are more aligned with the line-of-sight (e.g., $|\cos\theta|>0.8$).
Figure 7: This plot shows the distribution of the $L_{X}$ extracted from the halos in the triaxial XSB map (‘tri’, solid lines), in the XSB map with spherical halos plus intrinsic scatter (‘sph+var’, dashed lines), and XSB map with triaxial halos with intrinsic scatter (‘tri+var’, dotted lines), to that extracted from the spherical BP XSB map (‘sph’), in different halo mass bins. It shows that the scatter due to triaxiality is smaller than the intrinsic scatter. Lower mass halos have larger intrinsic scatter and larger triaxial scatter, compared to higher mass halos.
@@ 4.1 Extrinsic Scatter in X-ray Luminosity–Halo Mass Scaling Relation
Excluding halo core regions can significantly reduce intrinsic scatter in X-ray luminosity. By omitting pixels within a circular aperture of $0.2R_{500c}$ around each halo’s center, the intrinsic scatter decreases from $29\%$ to $22\%$ for groups and from $20\%$ to $11\%$ for massive clusters. However, the scatter due to triaxiality remains unchanged after core excision, maintaining approximately $12\%$ at the group scale and $7\%$ at the cluster scale, which constitutes nearly half of the total scatter. These results indicate that the halo outskirts are the primary contributors to triaxial scatter. Our results are qualitatively similar to the previous work based on hydrodynamical cosmological simulations (e.g., Battaglia et al., 2012a).
@@ 4.2 Bias in X-ray and Thermal SZ Power Spectra due to Intrinsic Scatter in ICM Profiles
Figure 8: Plots showing the effect of intrinsic scatter of pressure and density profiles on the X-ray and tSZ auto- and cross-angular power spectra. The left, middle, and right panels show the X-ray auto-power spectra, X-ray/tSZ cross-power spectra, and tSZ auto-power spectra, measured from the Uchuu BP maps respectively. In each panel, the blue and orange lines represent the power spectra with and without intrinsic scatter in the halo profiles. The bottom sub-panels show the ratios of the power spectra with intrinsic scatter in the halo profiles, to those without. Including intrinsic scatter in the gas profiles increases the normalizations of the power spectra. Note that the errorbars in each of the power spectrum are much smaller than the difference between the spectra. Also note that the tSZ auto-power spectrum is unit-less.
@@
The angular power spectrum of clusters is dominated by the one-halo term $C^{\rm 1h}$ at most scales of interest ($\ell>10$), so we will ignore the two-halo term. The observables ${\cal A}^{\prime}$ and ${\cal B}^{\prime}$ for a random halo in a given mass and redshift bin are expected to deviate from their expectation values ${\cal A}$ and ${\cal B}$ as,
${\cal A}^{\prime}$ $=$ ${\cal A}+\delta{\cal A},$ (33)
${\cal B}^{\prime}$ $=$ ${\cal B}+\delta{\cal B},$
where $\delta{\cal A}$ and $\delta{\cal B}$ are random variables. In the simplest halo model, where the $\delta$ terms are zero, ${\cal A}^{\prime}={\cal A}$ and ${\cal B}^{\prime}={\cal B}$, and resulting $C^{\rm 1h}$ term is the same as in Equation (4.2). However, in a more general scenario where the $\delta$ terms are non-zero, the $C^{\rm 1h}$ term becomes,
@@
×$\times\left({\cal A}^{*}_{\ell}(M,z){\cal B}_{\ell}(M,z)+\langle\delta{\cal A}^{*}_{\ell}(M,z)\delta{\cal B}_{\ell}(M,z)\rangle\right),$
[…]
Interpretation of the power spectrum based on the halo model usually overlooks the intrinsic scatter in the halo gas profiles. As a result, the halo model approach typically underestimates the actual amplitude of the power spectrum. Therefore, the intrinsic scatter profile provides an explanation for the differences in the power spectrum between simulations and analytic halo models (e.g., Battaglia et al., 2012b). Thus, cosmological parameters, such as $S_{8}$, inferred by statistical inference based on the halo model, are often overestimated to account for the additional power from intrinsic scatter.